# AC-LIO: Towards Asymptotic and Consistent Convergence in LiDAR-Inertial Odometry

### v1

@@ Abstract
Existing LiDAR-Inertial Odometry (LIO) frameworks typically utilize prior state trajectories derived from IMU integration to compensate for the motion distortion within LiDAR frames, and demonstrate outstanding accuracy and stability in regular low-speed and smooth scenes. However, in high-speed or intense motion scenarios, the residual distortion may increase due to the limitation of IMU’s accuracy and frequency, which will degrade the consistency between the LiDAR frame with its represented geometric environment, leading pointcloud registration to fall into local optima and consequently increasing the drift in long-time and large-scale localization.
To address the issue, we propose a novel asymptotically and consistently converging LIO framework called AC-LIO. First, during the iterative state estimation, we backwards propagate the update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration. Second, considering the weak correlation between the initial error and motion distortion of current frame, we propose a convergence criteria based on pointcloud constraints to control the back propagation. The approach of guiding the asymptotic distortion compensation based on convergence criteria can promote the consistent convergence of pointcloud registration and increase the accuracy and robustness of LIO. Experiments show that our AC-LIO framework, compared to other state-of-the-art frameworks, effectively promotes consistent convergence in state estimation and further improves the accuracy of long-time and large-scale localization and mapping.
@@ I INTRODUCTION
Figure 1: The basic principle of AC-LIO framework. Conventional LIO typically consider initial distortion compensation via IMU integration, yet the residual distortion may prevent the registration from further convergence. In contrast, our AC-LIO employs an asymptotic backpropagation strategy to promote consistent convergence in state estimation.
@@
With the development of Simultaneous Localization and Mapping (SLAM) technology, the perception of robots on the external environment and their own motion is becoming more and more precise [1, 2]. As a key branch, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving, unmanned logistics and other fields, for its high robustness and accuracy [3, 4]. Compared to Visual-Inertial Odometry (VIO), the LIO system can directly obtain depth observations of surroundings [5], without the need to estimate the depth of feature points [6, 7], and exhibits better adaptability to different lighting conditions [8, 9].
@@
However, as LIO system continues to evolve, there still remain several fundamental issues to be addressed [10, 11], one of which is the conflict between continuous motion and rigid pointcloud registration [12]. Since the acquisition of LiDAR stream is often accompanied with the ego-motion of vehicle, the laser points with different timestamps actually lie in different coordinate frames.
Existing frameworks typically only consider using IMU integration for the initial compensation of motion distortion [13, 14, 15], without further optimization of the residual distortion in state estimation. But in high-speed or violent motion scenarios, the residual motion distortion tends to increase due to the limitations of IMU’s accuracy and frequency, which would suppress further convergence of pointcloud registration and decrease odometry accuracy in long-term, large-scale scenarios [16].
@@
To address this, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework. First, during the iterative state estimation, we backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration. This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of state estimation and improving the accuracy of long-time and large-scene localization. Compared to other strategies such as continuous-time estimation or frame segmentation, our method can be seamlessly embedded into the iterative Kalman Filter framework [17], offering great flexibility and scalability.
@@
Second, considering the weak correlation between the initial error and motion distortion of current frame, we also propose a convergence criteria based on the point-plane residuals to control the back propagation. This strategy ensures, We use this criteria to evaluate the convergence of the pointcloud registration, ensuring the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, and thus guaranteeing the robustness of LIO system.
@@
  • We implement a novel asymptotically and consistently converging LIO framework. During the iterative update, we back propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration, promoting the consistent convergence of registration and improving the accuracy in long-time and large-scale scenarios.
  • We introduce a convergence criteria based on point-plane residuals to control the back propagation. We utilize the criteria to evaluate the convergence of registration, ensuring the back propagation to be performed only upon sufficient convergence of the previous registration, thus further enhancing the system robustness.
  • Our method can be easily and seamlessly integrated into traditional iterative Kalman Filter framework. Compared to other optimization strategies for motion distortion, our framework introduce no extra estimates or states, making the asymptotic compensation less computationally expensive but highly flexible and scalable.
  • We conduct extensive experiments on AC-LIO in a series of real-world datasets with various scenarios. The results show that our proposed AC-LIO exhibits higher accuracy in long-term localization compared to the state-of-the-art open-source methods.
@@ II RELATED WORK
Since the motion distortion can degrade the consistency between the pointcloud and the geometric environment, and thus suppress the sufficient convergence of registration. To mitigate the adverse effect of motion distortion, existing LiDAR(-inertial) odometry adopt various strategies to appropriately compensate for this distortion.
@@
For LiDAR-only odometry, most studies utilize the constant velocity model to compensate for motion aberrations. Studies such as LOAM [18] and Lego-LOAM [19] have achieved significant effects with the assumption of constant velocity motion during the corresponding time of LiDAR frames. However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth. In case of drastic velocity variation, the constant velocity model cannot accurately characterize the current motion state, and this will result in increased drift and even odometry divergence.
@@
As for LiDAR-inertial odometry, the IMU observation can provide relatively reliable prior state information. This provides a new scheme for compensating motion distortion via IMU integration. FAST-LIO [20] employs the IMU state backpropagation to correct the distorted pointcloud, and then conducts state estimation through Error State Kalman Filter. LIO-SAM [21] utilizes the IMU pre-integration to de-skew pointcloud, and combines multiple constraints including LiDAR, IMU, and GPS through the factor graph. Although these methods can mitigate the impact of pointcloud distortion to some extent, there are still some limitations without considering the further optimization of residual motion distortion.
To characterize trajectorie more precisely in state estimation, many frameworks come up with new optimization solutions. As pure LiDAR framework, I2EKF-LO [22] utilizes additional iteration to specifically estimate motion distortion, this approach can suppress distortion but the extra iteration may sacrifice computational efficiency. CLINS [23] introduces the B-spline model [24] for continuous time state estimation. However, the unconstrained trajectory would not always conform to the B-spline model due to time resolution limitations in high-speed, dynamic motion. Moreover, the continuous-time optimization strategy may introduce a higher computational cost. Point-LIO [25] adopts a point-by-point update strategy to avoid the motion distortion issue and realize extremely high-frequency odometry. Compared to frame based registration, the point-by-point strategy is more likely to degenerate in adverse scenarios. FR-LIO [26] divides the LiDAR frame into several sub-frames based on IMU observation. This approach can increase the system update frequency, thus reducing the impact of motion distortion. Besides, FR-LIO strengthens the constraints between subframes via RTS smoother [27], but still faces the risk of subframe degradation.
@@ III-A System Overview
Figure 2: The system overview of our AC-LIO framework. During iterative ESKF, asymptotic compensation of pointcloud distortion will be conducted based on the convergence criteria, to promote consistent convergence of state estimation.
The system overview of our AC-LIO framework is shown in Fig. 2. IMU observations are fed into the forward propagation module for integration to obtain prior trajectories. Then, the LiDAR data utilize the prior state chain for initial compensation, next passed into the Error State Kalman Filter (ESKF) for state estimation [28], where the average point-to-plane residual (APR) is calculated after each update.
During the iteration, if the APR of previous LiDAR frame is lower than the convergence criteria and the current APR exceeds it, then the backpropagation of update residual will be executed. After backpropagation, the LiDAR frame needs to be re-compensated to promote the consistent convergence of state estimation. Otherwise, if the conditions are not met, the backpropagation will be skipped and iteration continues. Once reaching the maximum of iterations, the system will merge the LiDAR frame under the posterior pose into pointcloud map and publish the current odometry.
@@ III-B ESKF State Estimation
$\begin{split}\mathbf{x}&\doteq\begin{bmatrix}\mathbf{R}_{I}^{G}&\mathbf{p}_{I}^{G}&\mathbf{v}_{I}&\mathbf{b}_{\bm{\omega}}&\mathbf{b}_{\mathbf{a}}&\mathbf{g}^{G}\end{bmatrix}^{T}\\ \mathbf{i}&\doteq\begin{bmatrix}{\bm{\omega}}_{m}&{\mathbf{a}}_{m}\end{bmatrix}^{T}\\ \mathbf{n}&\doteq\begin{bmatrix}\mathbf{n}_{\bm{\omega}}&\mathbf{n}_{\mathbf{a}}&\mathbf{n}_{\mathbf{b}\bm{\omega}}&\mathbf{n}_{\mathbf{b}\mathbf{a}}\end{bmatrix}^{T}\end{split}$ (1)
Where $\mathit{G}$ and $\mathit{I}$ denote the global and the IMU frames, $\mathbf{R}_{I}^{G}$, $\mathbf{p}_{I}^{G}$ and $\mathbf{v}_{I}^{G}$ are IMU’s orientation, position and linear velocity in $\mathit{G}$, and $\mathbf{g}^{G}$ is the local gravity. The $\mathbf{n}_{\bm{\omega}}$, $\mathbf{n}_{\mathbf{a}}$ are noises of angular velocity ${\bm{\omega}}_{m}$ and linear acceleration ${\mathbf{a}}_{m}$, the $\mathbf{b}_{\bm{\omega}}$, $\mathbf{b}_{\mathbf{a}}$ are IMU bias with $\mathbf{n}_{\mathbf{b}\bm{\omega}}$ and $\mathbf{n}_{\mathbf{b}\mathbf{a}}$ as their noises in Gaussian-Markov model.
@@
Denote $\mathbf{T}_{L}^{I}$ is the extrinsic between LiDAR and IMU, then the pointcloud can be transformed to frame $\mathit{I}$ via the prior pose chain. For a point $\mathbf{p}$ in LiDAR frame, its neighboring cluster can be searched by KNN in pointcloud map. A plane is then fitted to this cluster and we obtain the normal vector $\mathbf{u}$. Note one point of the cluster as $\mathbf{q}$ and $(\cdot)_{\wedge}$ as the upper triangular matrix of vector, then the point-plane residual $\mathbf{z}$ and measurement matrix $\mathbf{H}$ are as follows:
@@ III-C Asymptotic Distortion Compensation
In smooth motion scenarios, an initial compensation of motion distortion via IMU integration can basically satisfy the accuracy demand. However, when encountering violent motion scenarios, the residual motion distortion may increase due to the limitation of IMU’s frequency and accuracy. Compared to the noise of LiDAR observations, motion distortion may introduce bias in pointcloud registration, making it difficult to utilize the statistical properties of pointcloud to mitigate the adverse effects on state estimation.
With reference to the fixed interval smooth strategy, we selectively back propagate the update term $\delta\mathbf{x}$ based on the prior state chain during iteration. The noise between adjacent states can be expressed as follows:
$\begin{gathered}\delta\check{\mathbf{x}}_{i}\sim\mathcal{N}(\mathbf{0}_{18\times 1},\check{\mathbf{P}}_{i})\\ \delta\hat{\mathbf{x}}_{i+1}\sim\mathcal{N}(\mathbf{0}_{18\times 1},\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T})\end{gathered}$ (6)
Figure 3: The on-manifold propagation. The noise of current state involves the process noise and the propagation of previous state, the update term can be recursive backward according to this relationship.
During the on-manifold propagation of state $\mathbf{x}$, the noise of current error state $\delta\hat{\mathbf{x}}_{i+1}$ consists of the propagation of previous state $\delta\hat{\mathbf{x}}_{i}$ and the accumulated process noise. According to RTS smooth strategy [27], the contribution of previous state in current update term just corresponds to the proportion of the propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$. Consequently, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
@@
Furthermore, since there may exist numerous IMU states to maintain within the timeperiod of LiDAR frame, a sequential backward recursion can characterize the smoothed trajectory precisely. However, from an overall perspective, this treatment will increase much computational cost in iterative ESKF, but cannot bring additional accuracy improvement in subsequent asymptotic distortion compensation. Compared to state-by-state backpropagation, selectively updating several anchors in the prior state chain can further reduce the computational cost with little accuracy loss. To this end, we extend the backwards recursion to any states within the LiDAR frame, thereby reducing the computational complexity of backpropagation and distortion compensation.

### v2

@@ Abstract
Existing LiDAR-Inertial Odometry (LIO) methods typically utilize the prior state trajectory derived from the IMU integration to compensate for the motion distortion within LiDAR frames. However, discrepancies between the prior and actual trajectory can lead to residual distortions that compromise the consistency of the LiDAR frame with its corresponding geometric environment. This imbalance may result in pointcloud registration becoming trapped in local optima, thereby exacerbating drift during long-term and large-scale localization. To address the issue, we propose a novel asymptotically and consistently converging LIO framework dubbed AC-LIO. Our key idea is to back propagate current update term based on the prior state chain, and asymptotically compensate for the residual distortion during iteration. Moreover, considering the weak correlation between previous error and current distortion, we establish convergence criteria based on the pointcloud constraints to regulate the backpropagation. This method of guiding asymptotic distortion compensation using convergence criteria subtly enhances the consistent convergence of pointcloud registration, futher improving the accuracy and robustness of LIO system. Extensive experiments demonstrate that our AC-LIO framework significantly promotes consistent convergence in state estimation compared to prior arts, with about 30.4% reduction in average RMSE over the second best result, leading to marked improvements in the accuracy of long-term and large-scale localization and mapping.
@@ I INTRODUCTION
With the advancements in Simultaneous Localization and Mapping (SLAM) technology, robots are increasingly achieving precise perception of their external environments and their own motion[1, 2, 3]. As a vital subset of SLAM, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving [4], unmanned logistics [5], humanoid robot [6] and other fields, for its high robustness and accuracy [7, 8]. Compared to Visual-Inertial Odometry, the LIO system can directly obtain depth observations of surroundings [9], without the need to estimate the depth of feature points [10, 11], and exhibits better adaptability to different lighting conditions [12, 13].
Figure 1: (a) A principal comparison of our AC-LIO with traditional approaches. Conventional LIO typically considers an initial distortion compensation via IMU, yet the residual distortion may prevent the registration from further convergence. In contrast, our AC-LIO employs an asymptotic backpropagation to promote consistent convergence in state estimation. (b) Large scale localization and mapping result of BotanicGarden dataset. We achieve better accuracy performance than other benchmark frameworks on its sequence 1008_01 with VLP16 LiDAR.
@@
However, as LIO system continues to evolve, there still remain several fundamental issues to be addressed [14, 15, 16]. One of them is the conflict between continuous motion and rigid pointcloud registration [17]. Since the acquisition of LiDAR frames is often accompanied with the ego-motion of the vehicle, the LiDAR points with different timestamps actually lie in different coordinate systems.
As shown in Fig. 1, traditional methods [18, 19, 20] typically only consider using IMU integration for the initial compensation of motion distortion, without further optimization of the residual distortion in the state estimation. This residual distortion will compromise the consistency between the LiDAR frame with its corresponding geometric environment, and it grows with increasing motion non-linearity. Therefore, in conventional LIO systems, such residual error may suppress further convergence of pointcloud registration during the iterative update, ultimately impacting the odometry accuracy and causing severe drift.
@@
In this letter, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework. Our AC-LIO enjoys two key technical contributions. First, during the iterative state estimation, we introduce an asymptotic distortion compensation (Sec. III-C) method that backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration . This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of pointcloud registration and improving the accuracy of long-term and large-scene localization.
@@
Second, considering the weak correlation between initial error and motion distortion of current frame, we also propose the convergence criteria (Sec. III-D) based on the point-to-plane residuals to control the back propagation. The criteria is used to assess the convergence of the pointcloud registration and ensure that the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, hence guaranteeing the robustness of LIO system. Compared to other strategies such as continuous-time estimation or frame segmentation, our method leverages this criteria to construct a feedback mechanism that promotes state convergence without additional estimation, achieving a balance between localization accuracy and computational efficiency (Sec. IV-C). It can be seamlessly embedded into the iterative Kalman Filter framework [21], offering great flexibility and scalability.
@@
  • We propose a novel framework AC-LIO with asymptotic distortion compensation that showing plausible consistent convergence in long-term and large-scale scenarios.
  • We introduce the convergence criteria based on point-to-plane residuals to evaluate the convergence and guide back propagation. Back propagation is only performed upon sufficient convergence of previous registration, thus further enhancing the system robustness.
  • Compared to other optimization strategies for motion distortion, our framework introduce no extra estimates or states, making the asymptotic compensation less computationally expensive but highly flexible.
  • We conduct extensive experiments on AC-LIO in a series of real-world datasets with various scenarios. The results show that our method achieves higher accuracy compared to current state-of-the-art frameworks.
@@ II RELATED WORK
LiDAR-Only Odometry. Studies such as LOAM [22] and F-LOAM [23] has achieved significant performance with the assumption of constant velocity motion during the corresponding time of LiDAR frames. However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth. In case of drastic velocity variation, the constant velocity model cannot accurately characterize the current motion state, and this will result in increased drift and even odometry divergence. CT-ICP [24] optimize the start and end states of LiDAR frame, and construct a continuous time trajectory by interpolation. But the real movement may not follow the interpolated trajectory, as well as the logical constraint across adjacent frames is hard to determine. I2EKF-LO [25] utilizes extra prediction iteration to estimate motion distortion, yet the update of the process noise still depends on the hyperparameters. Traj-LO [26] leverages the high frequency of LiDAR points and employs more linear segments within frame. This approach allows more accurate trajectory characterization and suppression of motion aberrations during high speed motion, but also introduces more computational cost.
@@
Discrete State LiDAR-Inertial Odometry. For LiDAR-inertial odometry, the IMU observation can provide relatively reliable prior state information. FAST-LIO2 [27] employs the IMU state backpropagation to correct the distorted pointcloud, and then conducts state estimation through Error State Kalman Filter. LIO-SAM [28] utilizes the IMU pre-integration to de-skew pointcloud, and fuses multiple constraints including LiDAR, IMU, and GPS through the factor graph. In the subsequent LIO frameworks, SR-LIO [29] adopts a LiDAR frame reconstruction strategy, while AS-LIO [30] designs an adaptive sliding window guided by spatial overlap degree, both of which can boost the system update frequency and lower the impact of motion distortion. Semi-Elastic-LIO [31] replaces the interpolation model with IMU motion constraints on the basis of CT-ICP framework, thereby improving the stability of the inter-frame logical constraints. Although these methods can mitigate the impact of pointcloud distortion to some extent, there are still some limitations without considering the further optimization of residual motion distortion. Point-LIO [32] adopts a point-by-point update strategy to avoid the motion distortion issue and realize extremely high-frequency odometry. Compared to frame based registration, the point-by-point strategy is more likely to affected by dynamic objects and degenerate in adverse scenarios.
Continuous State LiDAR-Inertial Odometry. To characterize trajectorie more precisely in state estimation, many frameworks come up with new optimization solutions. CLINS [33] introduces the B-spline [34] for continuous time state estimation, while SLICT [35] employs continuous-time factors to represent LiDAR constraints within a sliding window. These frameworks further optimize motion distortion with continuous-time model. However, the unconstrained trajectory would not always conform to the predefined model due to time resolution limitations in high-speed, dynamic motion. Moreover, the continuous-time optimization strategy may introduce a higher computational cost. DLIO [36] uses analytical equations to construct a continuous trajectory for precise motion correction, after coarse deskewing through IMU integration. But the method remains limited by IMU’s inherent accuracy and neglects further undistortion during the registration process.
@@ III METHODOLOGY
TABLE I: Important Notations
Symbols Definition
[…]
$I,L,G$ The IMU, LiDAR, and global coordinate frame.
$\mathbf{x}^{gt}$ The ground truth of state $\mathbf{x}$.
$\mathbf{x}_{i}$ The vector of state $\mathbf{x}$ at timestamp $i$.
$\hat{\mathbf{x}}$ The priori estimation of state $\mathbf{x}$.
$\hat{\mathbf{x}}^{j},\delta\hat{\mathbf{x}}^{j}$ The $j$-th update of state $\mathbf{x}$ and error state $\delta\mathbf{x}$.
$\bar{\mathbf{x}},\check{\mathbf{x}}$ The updated state $\mathbf{x}$ with or without back propagation.
$\mathbf{F_{a}}^{i|i+n}$ The propagation matrix from state $i$ to $i+n$.
$\mathbf{G}_{i}^{i+n}$ The backward gain matrix of state $i$ with respect to $i+n$.
$\mathbf{z},z$ The vector and its scalar value of point-to-plane residual.
$\phi$ The incident angle of single laser beam on the fitted plane.
$\sigma_{t}$ The expectation of average point-to-plane residual.
@@ III-A System Overview
Figure 2: (a) The system overview of our AC-LIO framework. During iterative ESKF, asymptotic compensation of pointcloud distortion is conducted based on the convergence criteria, to promote consistent convergence of state estimation. (b) The on-manifold propagation. The blue shows the forward propagation of IMU integration. The green indicates the backward recursion of the update term according to the propagation matrix and noise relation.
The system overview of our AC-LIO framework is shown in Fig. 2, and the important notations is shown in Table I. First, the IMU observations are fed into the forward propagation module to obtain prior trajectories. Then the LiDAR data utilize the prior state chain for initial compensation and then sent to the Error State Kalman Filter (ESKF) for state estimation [37], where the average point-to-plane residual (APR) is calculated after each update.
During the iteration, if the APR of previous LiDAR frame is lower than the convergence criteria and the current APR exceeds it, then the backpropagation of update residual will be executed. The LiDAR frame will be re-compensated to promote the convergence of registration after each backpropagation. Otherwise, if the conditions are not met, the backpropagation will be skipped and iteration continues. Once reaching the maximum of iterations, the system will merge the LiDAR frame under the posterior pose into pointcloud map and publish the current odometry.
@@ III-B ESKF State Estimation
$\begin{split}\mathbf{x}&\doteq\begin{bmatrix}\mathbf{R}_{I}^{G}&\mathbf{p}_{I}^{G}&\mathbf{v}_{I}&\mathbf{b}_{\bm{\omega}}&\mathbf{b}_{\mathbf{a}}&\mathbf{g}^{G}\end{bmatrix}^{T}\\ \mathbf{i}&\doteq\begin{bmatrix}{\bm{\omega}}_{m}&{\mathbf{a}}_{m}\end{bmatrix}^{T},~{}~{}\mathbf{n}\doteq\begin{bmatrix}\mathbf{n}_{\bm{\omega}}&\mathbf{n}_{\mathbf{a}}&\mathbf{n}_{\mathbf{b}\bm{\omega}}&\mathbf{n}_{\mathbf{b}\mathbf{a}}\end{bmatrix}^{T}\end{split}$ (1)
@@
Assume the extrinsic $\mathbf{T}_{L}^{I}$ between LiDAR and IMU is already known. For a point $\mathbf{p}$ in LiDAR frame, its neighboring cluster can be searched by KNN in pointcloud map. If a plane $S$ can be fitted to this cluster, then we will obtain its normal vector $\mathbf{u}$. Note one point of the cluster as $\mathbf{q}$ and $(\cdot)_{\wedge}$ is the upper triangular matrix of vector, then the point-to-plane residual $\mathbf{z}$ and measurement matrix $\mathbf{H}$ are as follows:
@@ III-C Asymptotic Distortion Compensation
Compared to the noise of LiDAR observations, motion distortion may introduce bias in pointcloud registration, making it difficult to utilize the statistical properties of pointcloud to mitigate the adverse effects on state estimation.
During the on-manifold propagation of state $\mathbf{x}$, the noise of current error state $\delta\hat{\mathbf{x}}_{i+1}$ consists of the propagation of previous state $\delta\hat{\mathbf{x}}_{i}$ and the accumulated process noise. With reference to the Rauch-Tung-Striebel fixed interval smooth strategy [39], we can back propagate the update term $\delta\mathbf{x}$ based on the prior state chain. The contribution of previous state in current update term just corresponds to the proportion of propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$. Then, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
@@
Furthermore, this sequential backpropagation method can be extend to any states within the LiDAR frame. It allows us to selectively update several anchor in the prior state chain, thus reducing the computational complexity of backpropagation and distortion compensation. As for any states $\mathbf{x_{i}}$ and $\mathbf{x_{i+n}}$ in the prior state chain, the propagation matrix and backward gain matrix can be noted as $\mathbf{F_{a}}^{i|i+n}$ and $\mathbf{G}_{i}^{i+n}$, then the backpropagation relationship should be: